[Brahms-l] Shift report 20050205 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 05 2005 - 16:04:01 EST
1. good store when I arrived at CH. Later they got vacuum problem and Radiation Monitor Failure twice. Finaly, we got very good beam around 2:15pm, but beam was lost at 3:40pm for vacuum problem.
2. Runs for 12A861 + 45A2255
 RunNo    Trig6    Trig3   
12806     182k      187k
12807     149k      151k
12808     159k      179k 
12809      6k        12k  (beam dumped)
TPC pedestal run 12810
Non-TPC pedestal run 12811
change polarity of D5 from A->B
12A861 + 45B2255
RunNo     Trig6    Trig3
12812     22k      21k
12813     162k     168k
12814     97k      110k (beam dumped)

3. None found



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 5 16:04:23 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 05 2005 - 16:04:45 EST